OpenStudy (anonymous):

Mass and Percent Abundance of Different Isotopes of Element A Isotope Mass of the Isotope (amu) % Abundance A-32 31.6800 83.04 A-33 32.6700 0.5000 A-34 33.6600 0.07000 A-37 36.6300 16.39

OpenStudy (anonymous):

@Abmon98

OpenStudy (somy):

Multiply each % to its mass

OpenStudy (somy):

then add them up

OpenStudy (somy):

afterwards divide the result by 100

OpenStudy (somy):

\((31.6800 \times 83.04)+(32.6700 \times 0.5000)+\) \((33.6600\times 0.07000)+(36.6300 \times 16.39)\) = get the answer your answer/100= the final answer you want